OpenStudy (preetha):

How many OpenStudiers visited since the dawn of time? Answer is in the newsletter - in your email.

OpenStudy (preetha):

Add contact@openstudy.com to your email address book so you receive these newsletter emails.

OpenStudy (amistre64):

42 ... im going with 42

OpenStudy (preetha):

Haha. I wonder how many others get that reference!

OpenStudy (astrophysics):

*Raises hand* Huge Douglas Adams fan haha, so with that said I agree with amistre!

OpenStudy (e.mccormick):

He beat me to it. =(

OpenStudy (compassionate):

42

OpenStudy (e.mccormick):

@Compassionate weak.

OpenStudy (e.mccormick):

\(\huge \dfrac{42^{42}}{x}\) where x is the # that makes it the correct answer.

OpenStudy (compassionate):

The only thing weak here is your mathematical reasoning ;)

OpenStudy (preetha):

I actually want everyone to go see the OS newsletter in their email please...(-:

OpenStudy (e.mccormick):

Well, we don't want to give it away! They need to research the answer and learn! So they need to look at the email. Or solve: \(\huge \left(\ln 42^x + x^{42^e}\right)\cdot \pi^{42i}\) Where the answer is something in the email. \(\ddot{\dot\smile}\)

OpenStudy (skyz):

42, The meaning of life :)

OpenStudy (skyz):

I saw it preetha :3